A small fraction of rope hangs from a frictionless table,if the length of rope is L, find the speed with which the rope will leave the table

Grade:12th pass

1 Answers

Kartik
7 years ago
Let small part of chain hang be x and let this chain with small part move by dx so its potential energy is Mxg×dx/L integrate to give Mgx^2/2L .Put limit from 0 to L and potential energy is MgL/2. So use energy conversation to get Mv^2/2=MgL/2 and v=(gl)^1/2